### v1

# Enhancement of the molecular electron chirality by electronic excitation

Abstract
Previously, it was predicted that the parity-violating energy difference (PVED) between enantiomers and the molecular electron chirality (MEC), which is the integrated value of the electron chirality density over the entire molecule are enhanced for ionization or electronic excitation. Following the confirmation of the enhancement of the PVED by electronic excitation in our previous papers, the present work confirms that electronic excitation enhances the MEC for H2Te2 and H2Se2. The PVED in the first excited state is explained by the contribution from the highest occupied molecular orbital if the PVED contribution from the highest occupied molecular orbital in the ground state dominates over any other contributions. In this paper, it is checked whether this explanation can also be applicable to the MEC. It is also reported that the MEC is not the simple sum of contributions from each atom and the mixing by the superposition between atomic orbitals of different atoms is important.
I Introduction
It is well-known that some molecules have chiral structures, while the electron also has its chirality, i.e. the handedness of the electron can be defined. A free electron does not have a specific charge distribution and the handedness of the electron cannot be defined by the distribution. However, the electron has the spin, and the inner product of the spin and momentum can define one type of handedness quantity of the electron. This definition of the handedness is the helicity. When the external field is negligible, this helicity is a conserved quantity in time, since the helicity commutes with the Hamiltonian. However, the helicity has a different ratio of the electron handedness in a different inertial frame and is not Lorentz invariant definition. The chirality of the electron is defined as the same handedness ration in any inertial frame. This is a relativistic extension of the helicity as explained in the next section.
Nature chooses the chirality as the handedness quantity of electrons. First, the weak interaction, which is one of the four fundamental interactions in particle physics, distinguish left- and right-handed electrons by the chirality. For example, this interaction is mediated by W and Z gauge bosons, and the W boson has the coupling with left-handed electrons and not with right-handed ones. In another example of the distinction, the torque density for the electron spin in quantum field theory is predicted to be dependent on the electron chirality density [1, 2]. In this theoretical prediction, the local torque contribution, which is named the zeta force, is given as the gradient of the electron chirality density. Although this contribution of the torque density has not been observed in experiments, this is confirmed by numerical simulations [3].
Although the physical property of the electron is different in its chirality, free electrons lose their polarization in chirality sooner or later due to the chirality flip by the electron mass. For example, if electrons are initially polarized more in left-handed chirality, it attenuates to no polarization state. The left-handed component flips to the right-handed one more than the inverse process (from right-handed to left-handed), since the ratio of the left-handed component is larger than the right-handed one.
However, electrons in chiral molecules, which are parity-violating objects, are considered to be polarized in the chirality, since a state polarized in the electron chirality may be the most stable in the steady state due to the spin-orbit interaction. Even though the electron mass flips the electron chirality, the attenuation goal is the polarized state in the chirality.
One phenomenon related to the electron chirality is the parity-violating energy difference (PVED) between enantiomers. This energy difference is induced by the interaction of Z-boson exchange between nuclei and electrons and is proportional to the electron chirality density difference at nuclei between enantiomers. Although this energy difference has not been confirmed in observational experiments despite hard efforts [4], theoretical works with quantum chemistry computations confirm nonzero PVED in many chiral molecules [5, 6, 7, 8]. Nevertheless, this energy difference is related to only the electron chirality at nuclei. The electron chirality for the entire molecule of H2Te2 is reported to be nonzero [9]. Then, in Ref. [6, 10, 11], it is systematically shown that chiral molecules have generally a nonzero integrated value of electron chirality density over the entire molecule. The existence of this chirality is named Electron Chirality in Chiral Molecules (ECCM) [11].
The value of ECCM is not large, and phenomena related to ECCM are not so many. Nevertheless, some phenomena are studied by our group. One topic is the homochirality in nature [12]. In an enantiomeric pair, the number of the left-handed electron is different, though the total electron number is the same. Hence, one enantiomer has a larger reaction rate of weak interaction than the other one. As a result, one type of an enantiomeric pair is more lost by collisions with particles with weak charge in space or on the earth [6]. Another example is the Chirality-Induced Spin Selectivity (CISS) [13]. CISS is the strong spin polarization when electrons pass through chiral organic molecules. It has not been clarified what makes large spin polarization. For this puzzle, the study from the viewpoint of the torque may be a key to the solution instead of the spin-orbit interaction, which is in the concept of energy. The torque for the electron spin is dependent on the electron chirality density, and the torques induced by the left- and right-handed electrons are opposite. ECCM predicts opposite electron chirality distribution in enantiomeric pair molecules. The zeta force may be the driving torque for the CISS. The correlation between the spin polarization of the CISS and the magnitude of the zeta force was confirmed for alanine and helicene molecules [14].
[…]
In the present paper, it is confirmed that the enhancement by electronic excitation also occurs for the MEC. As target molecules, H2Te2 and H2Se2 are chosen. In Refs. [7, 8], H2O2 and H2S2 are also studied for the enhancement of the PVED. For the MEC, the computation of molecules including only light elements with small spin-orbit interaction such as H2O2 and H2S2 requires very large basis sets and high cost [6, 11]. Hence, H2O2 and H2S2 are omitted in the present study for the excited states. However, we consider that H2O2 and H2S2 will show results similar to H2Te2 and H2Se2 due to the similar patterns of the total chirality in the ground state [6, 11] and the enhancement of the PVED [7, 8]. After the enhancement of the MEC is confirmed, it is studied whether the CBE hypothesis can also be applicable to the enhancement of the MEC.
This paper is organized as follows. In the next section, the chirality of the electron is introduced, and then the computational method and details are explained in Sec. III. In Sec. IV, our results of the enhancement of the MEC are shown. First, the enhancement of the MEC by electronic excitation is confirmed and then the CBE hypothesis is applicable to the enhancement of the MEC. The last section is devoted to the summary.
II Theory
The MEC, the integrated value of the electron chirality density over the entire molecule, is defined by the following formula,
$\int d^{3}x\psi^{\dagger}(x)\gamma_{5}\psi(x),$ (1)
where $\gamma_{5}\equiv i\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3}$ with the gamma matrices, $\gamma^{\mu}~{}(\mu=0-3)$ and $\psi$ is the electron wave function in a four-component representation. The electron chirality density, $\psi^{\dagger}(x)\gamma_{5}\psi(x)$, can be reduced to the combination form of the left- and right-handed electron components,
$\psi^{\dagger}(x)\gamma_{5}\psi(x)=\psi_{R}^{\dagger}(x)\psi_{R}(x)-\psi_{L}^{\dagger}(x)\psi_{L}(x).$ (2)
the left- and right-handed electron wave functions are defined as
$\psi_{R,L}(x)=P_{R,L}\psi(x),$ (3)
with the projection operators,
$P_{R}=\frac{1+\gamma_{5}}{2},~{}~{}~{}P_{L}=\frac{1-\gamma_{5}}{2}.$ (4)
The Lorentz transformation commutes with $\gamma_{5}$ [20] and the chirality i.e. the ratio between the left- and right-handed electrons are not changed in a different inertial frame. From $\gamma_{5}$ operator, it is seen that the chirality is a relativistic extension of the helicity. This operator is cast into another form, which is the product of the velocity and spin operators,
$\gamma_{5}=\alpha_{x}\Sigma_{x}=\alpha_{y}\Sigma_{y}=\alpha_{z}\Sigma_{z}=\frac{1}{3}\sum_{i=1}^{3}\alpha^{i}\Sigma^{i}.$ (5)
Here, $\alpha^{i}=\gamma^{0}\gamma^{i}$, and $\Sigma^{i}=\gamma^{0}\gamma^{i}\gamma_{5}$, which is the $4\times 4$ extension of the Pauli matrices. Therefore, the latter operator is the spin operator. The velocity operator is known to be defined as $\hat{v}^{i}=c\alpha^{i}$ with $\alpha^{i}$ matrices [20]. Hence, $\gamma_{5}$ operator is the product of the velocity (divided by $c$) and spin operators in relativistic quantum theory. However, the observable of the chirality density is not the product of the observed velocity and spin, $\psi^{\dagger}c\alpha^{i}\psi\psi^{\dagger}\Sigma^{i}\psi$, and the observable of the chirality density is $\psi^{\dagger}\gamma_{5}\psi$. The former is not Lorentz invariant.
In this paper, the CBE hypothesis [8] is checked for the MEC. In this paragraph, the CBE hypothesis is reviewed. This hypothesis requires the following conditions: the PVED is derived as the result of the cancellation between contributions from the HOMO and other orbitals, the electron in the HOMO in the ground state is dominantly excited, and other occupied orbitals in the ground state are not affected by the excitation. These conditions are satisfied in the first electronic excitation state in many molecules [8]. If these conditions are satisfied, the PVED in the first excited state is much larger than the PVED in the ground state. In addition, if the HOMO contribution to the PVED dominates over any other contributions, the PVED value is roughly estimated only with the contribution from the HOMO,
$E_{\rm PV}({\rm CBE})=-E_{\rm PV,HOMO}.$ (6)
Here $E_{\rm PV,HOMO}$ is the contribution from one electron in the Kramers pair of the HOMO to $E_{\rm PV}$. (Two electrons in the same Kramers pair give the same contribution to the electron chirality.) As a more improved estimate, contributions from other orbitals are taken into account,
$E_{\rm PV}({\rm IE})=\sum_{i}^{\rm unocc}r_{i}^{U}E_{\rm PV,\it i}-\sum_{i}^{\rm occ}r_{i}^{O}E_{\rm PV,\it i},$ (7)
where $E_{\rm PV,\it i}$ is the contribution from one electron in the $i$-th occupied or unoccupied orbital, and $r_{i}^{O}$ ($r_{i}^{U}$) is the excitation ratio of the electron in the $i$-th occupied orbital (to the $i$-th unoccupied orbital). The former simple estimate is quite useful, since only Hartree-Fock (HF) or density functional theory computations are needed, while the improved estimate (IE) requires a computation beyond self-consistent field computations.
While this CBE hypothesis is about the enhancement of the PVED, the enhancement of the MEC is considered to occur in the same cancellation-breaking mechanism. For the CBE hypothesis about the MEC, the conditions are the same as the PVED. The estimate of the MEC is given as the HOMO contribution to the MEC, that is,
$-\int d^{3}x\psi_{\rm HOMO}^{\dagger}(x)\gamma_{5}\psi_{\rm HOMO}(x),$ (8)
where $\psi_{\rm HOMO}$ is the wave function of the HOMO. The improved estimate with other orbital contributions is given in the following expression,
$\sum_{i}^{\rm unocc}r_{i}^{U}\int d^{3}x\psi_{i}^{\dagger}(x)\gamma_{5}\psi_{i}(x)-\sum_{i}^{\rm occ}r_{i}^{O}\int d^{3}x\psi_{i}^{\dagger}(x)\gamma_{5}\psi_{i}(x),$ (9)
where $\psi_{i}$ is the wave function of one of the Kramers pair of the $i$-th orbital. In this paper, the CBE hypothesis is checked for the MEC.
III Computational detail
The enhancement of the PVED of H2Te2 and H2Se2 was confirmed in our previous works [7, 8]. In this work, the MEC of H2Te2 and H2Se2 in electronic excited states is studied with quantum chemistry computations and the enhancement of it is confirmed. Then the enhancement mechanism of the MEC is compared with that of the PVED by using these molecules. The MEC of these molecules in the ground state was studied in previous works [6, 11, 9].
The structures of these molecules are chosen to be the same as our previous work [8]. The difference in structure between ground and excited states is ignored for simplicity, since the purpose of this paper is the confirmation of the enhancement of the MEC by electronic excitation and to study the mechanism of the enhancement. The structures of H2Te2 and H2Se2 were determined by geometrical optimization computations with the DIRAC19 program package [21, 22]. These computations were performed with the HF method with Dirac-Coulomb Hamiltonian and the dyall.ae2z basis set [23, 24]. After the geometrical optimization, the dihedral angle $\phi$ is taken as a free parameter and the structure dependence of the enhancement is studied for the change of $\phi$.
Computations of relativistic four-component wave functions are carried out with DIRAC19. Computations of excited states are based on Equation-of-Motion Coupled-Cluster (EOM-CC) theory. The MEC in excited states is calculated by using the Finite-Field Perturbation Theory (FFPT) [25, 26] as explained later. These computations are also supported by the comparison with results by the Z-vector method of Coupled Cluster Singles and Doubles (CCSD). Electron chirality of a specific molecular orbital is calculated with the HF method. For HF computations, the Dirac-Coulomb-Gaunt Hamiltonian is used, and for CCSD and EOM-CCSD computations using the RELCCSD module [27, 28], molecular mean-field approximations to the Dirac-Coulomb-Gaunt Hamiltonian are used [29]. In all computations, the Gaussian nuclei density distribution is used [30].
In computations of H2Te2 and H2Se2, C2 symmetry is adopted, and the dyall.acv3z basis set [24, 31] is utilized. In CCSD and EOM-CCSD computations of H2Te2 and H2Se2, electrons in 4s4p4d (Te) and 3s3p3d (Se) orbitals are chosen for frozen core electrons. The upper limit cutoff of the active space is 70 Hartree (H2Te2) and 100 Hartree (H2Se2). The threshold of the convergence in EOM-CCSD computations is $10^{-10}$.
In our computations, the calculation of the MEC in excited states is done with the FFPT [25, 26]. The MEC is derived by the perturbation calculation with the perturbation Hamiltonian $H_{5}=\lambda\gamma_{5}$, where $\lambda$ is the perturbation parameter. The expectation value of this operator is
$\langle H_{5}\rangle=\lambda\int d^{3}x\psi^{\dagger}(x)\gamma_{5}\psi(x)=\lambda\int d^{3}x\langle\gamma_{5}\rangle.$ (10)
For this expectation value, the following relation is given with the Hellmann-Feynman theorem,
$\left.\frac{\partial E\left(\lambda\right)}{\partial\lambda}\right|_{\lambda=0}$ $=\left\langle\psi\left|\frac{\partial H}{\partial\lambda}\right|\psi\right\rangle=\int d^{3}x\psi^{\dagger}(x)\gamma_{5}\psi(x),$ (11)
where $H$ is the total Hamiltonian and $\langle\psi|$ and $|\psi\rangle$ are bra and ket. Using the finite difference method, $\partial E/\partial\lambda|_{\lambda=0}$ can be computed as
$\left.\frac{\partial E}{\partial\lambda}\right|_{\lambda=0}\simeq\frac{E(\lambda)-E(-\lambda)}{2\lambda}.$ (12)
This computation of $E(\pm\lambda)$ is done with the EOM-CCSD in the Dirac program. For the finite difference method, accurate computations require adequate value of $\lambda$. For too large $\lambda$, effects from second or higher-order derivatives cannot be neglected, while for too small $\lambda$, the effect of the perturbation is invisible, since the effect is buried in a numerical error. To search for an adequate value of $\lambda$, the values of the MEC in the ground states are compared with those derived with CCSD computations. Once the value of $\lambda$ is determined, the same value of $\lambda$ is used for excited states. In the present work, $\lambda=0.1$ is adopted for both H2Te2 and H2Se2 as explained later.
IV Results
In this section, the atomic units are used for the electron chirality, particularly, the MEC.
(a) H2Te2
Refer to caption(b) H2Se2
Figure 1: The molecular electron chirality in the ground and excited states as a function of the dihedral angle. The line named GS means the ground state. $nx$ lines means the $n$-th excited state obeying the $x$-symmetry.
Figure 1 shows the electron chirality in excited states of (a) H2Te2 and (b) H2Se2 as a function of the dihedral angle. The notation of excited states, $nx$, means the $n$-th excited state obeying the $x$-symmetry. The electron chirality in excited states have opposite sign and much larger values compared to that of the ground state (GS). The enhancement of the MEC as well as the PVED is confirmed. While the MEC in the ground state on the dihedral angle is a sigmoidal curve, in excited states a different pattern of the dependence is seen. All curves of the MEC have similar distributions. In their patterns, the MEC increases to 90 and flips the sign around $\phi=90^{\circ}$. Most excited states are derived as the excitation of one electron in the HOMO in the ground state. The excitation ratio of the HOMO is roughly 90% except for 2a and 2b states around $\phi=90^{\circ}$ for both molecules, where the ratio is 20-70%. The energies of these excited states were reported in Ref. [8]. For both molecules, 1a, 1b, and 2a ($\phi<90^{\circ}$) or 1a, 1b, and 2b ($\phi>90^{\circ}$) are in the same approximate triplet state. The rest 2a (2b) state is the singlet state for $\phi>105^{\circ}$ $(\phi<75^{\circ})$, and forms approximate triplet state with 3a and 3b for $90^{\circ}<\phi<105^{\circ}$ $(75^{\circ}<\phi<90^{\circ})$.
The most stable structure of these excited states of H2Te2 and H2Se2 are achiral ($\phi=0^{\circ}$ or $180^{\circ}$). Actually, the MEC for the optimized structure in these excited states is zero, indicating no enhancement for these molecules in these excited states. This study of H2Te2 and H2Se2 is not for the search for molecules with large MEC in excited states but for the confirmation of the enhancement in electronic excited states. Moreover, the optimized structure of electronic excited states 3a, 3b, and 4a have $\phi=90^{\circ}$ (3a and 3b) or $105^{\circ}$ (4a) [8]. Hence, some higher electronic excited states have chiral structures.
Table 1: Dependence of the MEC on $\lambda$ within the FFPT method in the electronic ground state of H2Te2 at $\phi=45^{\circ}$. Dev. represents the relative deviation of the FFPT result from the Z-vector one.
Method
$\lambda$ [a.u.]
$\int d^{3}x\langle\gamma_{5}(x)\rangle/10^{-5}$ [a.u.]
Dev. [%]
FFPT $1.0\times 10^{+1}$ $0.581$ $-64.2$
$1.0\times 10^{-0}$ $1.574$ $-2.8$
$1.0\times 10^{-1}$ $1.585$ $-2.2$
$1.0\times 10^{-2}$ $1.558$ $-3.9$
$1.0\times 10^{-3}$ $1.440$ $-11.1$
Z-vector - $1.621$
Table 2: Dependence of the MEC on $\lambda$ within the FFPT method in the electronic ground state of H2Se2 at $\phi=45^{\circ}$. Dev. represents the relative deviation of the FFPT result from the Z-vector one.
Method
$\lambda$ [a.u.]
$\int d^{3}x\langle\gamma_{5}(x)\rangle/10^{-6}$ [a.u.]
Dev. [%]
FFPT $1.0\times 10^{+1}$ $-0.323$ $-115.4$
$1.0\times 10^{-0}$ $1.998$ $-4.5$
$1.0\times 10^{-1}$ $2.025$ $-3.3$
$1.0\times 10^{-2}$ $1.984$ $-5.2$
$1.0\times 10^{-3}$ $1.985$ $-5.2$
Z-vector - $2.093$
As explained in the previous section, the usage of the adequate value of $\lambda$ is important for the evaluation of the MEC with the FFPT method. In our work, the value of $\lambda$ is determined by the comparison with the value of the MEC in the CCSD computation using the Z-vector method. The dependence of the value of the MEC on $\lambda$ in the FFPT computations is shown in Tables 2 and 2. For these computations, the dihedral angle is chosen to be $\phi=45^{\circ}$, since the integrated chirality is almost zero in optimized structure $(\phi=90^{\circ}$) and large at $\phi=45^{\circ}$. The deviation is the smallest at $\lambda=0.1$, and this value is adopted for both molecules.
Table 3: Dependence of the MEC on $\lambda$ within the FFPT method in excited states of H2Te2 at $\phi=45^{\circ}$.
$\lambda$ [a.u.]
State
$\int d^{3}x\langle\gamma_{5}(x)\rangle/10^{-5}$ [a.u.]
$1.0\times 10^{-0}$ 1a $-4.656$
1b $-2.182$
2a $-5.567$
2b $-6.325$
$1.0\times 10^{-1}$ 1a $-4.707$
1b $-2.205$
2a $-5.626$
2b $-6.389$
$1.0\times 10^{-2}$ 1a $-4.744$
1b $-2.240$
2a $-5.653$
2b $-6.425$
This single value of $\lambda$ was used for all excited states and all angles of the same molecule in the computation of Fig. 1. Even for excited states, the small change of $\lambda$ does not affect our results for excited states as seen in Table 3. The adequate value of $\lambda$ is determined by the ratio of the perturbation energy $\int d^{3}x\langle\gamma_{5}(x)\rangle$ to the molecular one. The change of the molecular energy is small for the excitation and the change of the dihedral angle compared to the increase of $\int d^{3}x\langle\gamma_{5}(x)\rangle$ by the excitation. The increase of $\int d^{3}x\langle\gamma_{5}(x)\rangle$ from the ground state at $\phi=45^{\circ}$ is several to ten times from Fig. 1. In Tables 2 and 2, the deviation is less than 3% (H2Te2) and 5% (H2Se2) for $\lambda=1.0$ and hence for ten times larger perturbation Hamiltonian $\lambda\int d^{3}x\langle\gamma_{5}(x)\rangle$ in the ground state at $\phi=45^{\circ}$ the deviation is confirmed to be acceptably small. Hence, the usage of the same $\lambda$ in excited states is considered to be appropriate.
(a) H2Te2
Refer to caption(b) H2Se2
Figure 2: Contributions from single molecular orbitals to the MEC as a function of the dihedral angle as well as the total electron chirality summed up all the occupied orbitals.
To understand the dependence of the MEC on the dihedral angle, contributions from single molecular orbitals are studied. In Fig. 2, contributions from HOMO, HOMO-1, and HOMO-2 are shown as a function of the dihedral angle. For both molecules, contributions from HOMO, HOMO-1, and HOMO-2 are much larger than the total electron chirality summed up all the occupied orbitals. These three contributions are almost canceled out. This is the key to the idea of the enhancement of the electron chirality and is consistent with the previous report [6]. The pattern of the HOMO contribution is opposite to that of the MEC of 1a and 1b excited states. This is consistent with the CBE hypothesis.
(a) H2Te2
Refer to caption(b) H2Se2
Figure 3: The prediction of the MEC by the CBE hypothesis.
In our previous paper [8], it is reported that the value of the PVED in the first excited state of some molecules can be predicted as the opposite sign of the HOMO contribution to the PVED in the CBE hypothesis. In this hypothesis, when in the ground state the PVED contribution from the HOMO is larger than the total value of occupied orbitals, the PVED in the first excited state is much larger than that in the ground state by breaking the cancellation between contributions from the HOMO and the other orbitals. Moreover, the PVED contribution from the HOMO in the ground state dominates over any contribution from other occupied orbitals, the PVED in the first excited state can be estimated from only the HOMO contribution. Since the MEC of H2Te2 and H2Se2 is also derived as a result of the cancellation between contributions from the HOMO and others orbitals [6], the conditions of the CBE hypothesis are satisfied for the MEC in H2Te2 and H2Se2.
In Fig. 3, the estimate of the MEC in the CBE hypothesis is shown as well as the improved estimate as introduced in Sec. II. The predictions by the CBE hypothesis well realize the patterns of the MEC in 1a and 1b states. Particularly, all predictions for $\phi>90^{\circ}$ are perfectly consistent with computational results of the MEC. For $\phi<90^{\circ}$ the prediction in the 1a state of H2Te2 realizes precisely the MEC curve, while prediction curves for other states have a few hundred percents deviations. This inconsistency is considered to arise from the modification of occupied orbitals from the ground state and the difference between HF and CCSD computations. Nevertheless, the patterns of curves are consistent with the computational results. The CBE hypothesis is considered to apply the MEC. At most points, the simple CBE estimate does not deviate from the prediction by IE. This is because the electron in the HOMO is dominantly excited at most points. The prediction by the improved estimate seems to be worse than the simple CBE estimate around $\phi=90^{\circ}$. This is because there are few data for IE points. If there were additional data around $\phi=90^{\circ}$, the IE prediction was better than Fig. 3.
(a) H2Te2
Refer to caption(b) H2Se2
Figure 4: Atomic contribution in the HOMO to the MEC as a function of the dihedral angle. For HOMO(Te, Se) and HOMO(H), the HOMO contribution to the MEC is divided into atomic contributions, where the contribution from atomic orbitals of only a single atom (Te, Se, or H) is taken into consideration.
Finally, the dependence of the MEC on the dihedral angle is discussed. It is known that the dependence of the MEC and PVED on the dihedral angle is almost the same in the ground state of H2Te2 and H2Se2 [6, 11, 9]. For excited states, MEC curves have different patterns compared to PVED ones. The pattern of MEC curves is like the teeth of a saw as shown in Fig. 1, while PVED curves have a narrow peak [8]. The difference in the dependence on the dihedral angle between the MEC and the PVED in the first excited states can be elucidated by the CBE hypothesis, and the dependence is almost explained by the HOMO contribution dependence on the dihedral angle as seen from Eqs. (6) and (8). Therefore, the difference in the dependence is the difference in the HOMO contribution. It is known that the PVED can be divided into contributions from each atom [17]. Hence, to understand the difference in the HOMO contribution more deeply, the HOMO contribution to the MEC is divided into atomic contributions, i.e. heavier (Te or Se) and lighter (H) atomic contributions. In Fig. 4, atomic contributions in the HOMO to the MEC is shown as a function of the dihedral angle. For lines of HOMO(Te, Se) and HOMO(H), the contribution from atomic orbitals of only a single atom (Te, Se, or H) is taken into consideration. These contributions are calculated with QEDynamics (module for Dirac) [32]. As seen from Fig. 4, the HOMO contribution to the MEC is not the simple sum of atomic contributions and the mixing by the superposition between atomic orbitals of heavier and lighter elements is inevitably important. This feature is a salient contrast to the PVED. In Ref. [9], it was implied that the MEC as well as the PVED is also the very atomic nature, since the electron chirality is derived as the product of the small and large components of the four-component wave function. Our result shows that the MEC is not so simple and not atomic nature. This is also the reason why computations of the MEC for molecules with only light elements such as H2O2 requires additional diffuse functions for basis sets, which reported in previous work [11].
V Summary
[…]
Acknowledgements.
This work was supported by Grants-in-Aid for Scientific Research (21H00072, and 22K12060). The work of N. K. was supported by Grant-in-Aid for JSPS Fellows (23KJ1188) and JST SPRING, Grant Number JPMJSP2110. In this research work we used the supercomputer of ACCMS, Kyoto University. The authors are grateful to Ayaki Sunaga for the contribution at the early stage of this work.

### v2

@@
 I Introduction
-It is well-known that some molecules have chiral structures, while the electron also has its chirality, i.e.
+It is well established that certain molecules exhibit chiral structures, while electrons also possess an intrinsic chirality, i.e.
 the handedness of the electron can be defined.
-A free electron does not have a specific charge distribution and the handedness of the electron cannot be defined by the distribution.
-However, the electron has the spin, and the inner product of the spin and momentum can define one type of handedness quantity of the electron.
-This definition of the handedness is the helicity.
-When the external field is negligible, this helicity is a conserved quantity in time, since the helicity commutes with the Hamiltonian.
-However, the helicity has a different ratio of the electron handedness in a different inertial frame and is not Lorentz invariant definition.
-The chirality of the electron is defined as the same handedness ration in any inertial frame.
-This is a relativistic extension of the helicity as explained in the next section.
-Nature chooses the chirality as the handedness quantity of electrons.
-First, the weak interaction, which is one of the four fundamental interactions in particle physics, distinguish left- and right-handed electrons by the chirality.
-For example, this interaction is mediated by W and Z gauge bosons, and the W boson has the coupling with left-handed electrons and not with right-handed ones.
-In another example of the distinction, the torque density for the electron spin in quantum field theory is predicted to be dependent on the electron chirality density [1, 2].
-In this theoretical prediction, the local torque contribution, which is named the zeta force, is given as the gradient of the electron chirality density.
-Although this contribution of the torque density has not been observed in experiments, this is confirmed by numerical simulations [3].
-Although the physical property of the electron is different in its chirality, free electrons lose their polarization in chirality sooner or later due to the chirality flip by the electron mass.
-For example, if electrons are initially polarized more in left-handed chirality, it attenuates to no polarization state.
-The left-handed component flips to the right-handed one more than the inverse process (from right-handed to left-handed), since the ratio of the left-handed component is larger than the right-handed one.
-However, electrons in chiral molecules, which are parity-violating objects, are considered to be polarized in the chirality, since a state polarized in the electron chirality may be the most stable in the steady state due to the spin-orbit interaction.
-Even though the electron mass flips the electron chirality, the attenuation goal is the polarized state in the chirality.
-One phenomenon related to the electron chirality is the parity-violating energy difference (PVED) between enantiomers.
-This energy difference is induced by the interaction of Z-boson exchange between nuclei and electrons and is proportional to the electron chirality density difference at nuclei between enantiomers.
-Although this energy difference has not been confirmed in observational experiments despite hard efforts [4], theoretical works with quantum chemistry computations confirm nonzero PVED in many chiral molecules [5, 6, 7, 8].
-Nevertheless, this energy difference is related to only the electron chirality at nuclei.
-The electron chirality for the entire molecule of H2Te2 is reported to be nonzero [9].
-Then, in Ref.
-[6, 10, 11], it is systematically shown that chiral molecules have generally a nonzero integrated value of electron chirality density over the entire molecule.
+A free electron does not have a specific charge distribution and the handedness of the electron should not be defined by the distribution.
+However, the electron possess spin, and the inner product of spin and momentum enables the definition of a specific handedness quantity known as helicity.
+In the absence of external fields, helicity is conserved over time, since it commutes with the Hamiltonian operator.
+It is noteworthy that helicity is frame-dependent and is not Lorentz invariant.
+In contrast, the chirality of an electron is defined as a quantity that maintains a consistent handedness ratio across all inertial frames.
+This concept represents a relativistic extension of helicity as elucidated in the next section.
+Nature inherently selects the chirality as the handedness quantity of electrons.
+Notably, the weak interaction, one of the four fundamental interactions in particle physics, differentiates between left- and right-handed electrons based on their chirality.
+For instance, this interaction is mediated by W and Z gauge bosons, and the W boson interacts with left-handed electrons but does not with right-handed electrons.
+Additionally, in the framework of quantum field theory, the torque density for the electron spin is predicted to be dependent on the electron chirality density [1, 2].
+In this theoretical prediction, the local torque contribution, which is named the zeta force, is expressed as the gradient of the electron chirality density.
+While this contribution to torque density has yet to be observed experimentally, it has been corroborated by numerical simulations [3].
+Although the chirality of an electron affects its physical properties, free electrons eventually lose their chirality polarization due to mass-induced chirality flips.
+For instance, if electrons are initially polarized predominantly in the left-handed chirality, this polarization will decay over time to a state of no net polarization.
+The left-handed component tends to convert to the right-handed component more frequently than the reverse process occurs, as the initial ratio of left-handed to right-handed components is greater.
+However, electrons within chiral molecules, which are parity-violating entities, are expected to exhibit polarization in chirality.
+This is because a state polarized in electron chirality may represent the most stable configuration in steady-state conditions due to spin-orbit interactions.
+Although the mass of the electron can induce chirality flips, the ultimate objective is to achieve a stable polarized state in chirality.
+One phenomenon associated with the electron chirality is the parity-violating energy difference (PVED) between enantiomers.
+This energy difference arises from the Z-boson exchange interaction between nuclei and electrons and is proportional to the difference in electron chirality density at the nuclei of the enantiomers.
+Although extensive experimental efforts have yet to confirm this energy difference [4], theoretical studies utilizing quantum chemistry computations indicate a nonzero PVED in many chiral molecules [5, 6, 7, 8].
+Importantly, this energy difference pertains solely to the electron chirality at the nuclei.
+The electron chirality for the entire molecule of H2Te2 was reported to be nonzero [9].
+Furthermore, systematic analyses in Ref.
+[6, 10, 11] demonstrate that chiral molecules generally exhibit a nonzero integrated value of electron chirality density over the entire molecular structure.
 The existence of this chirality is named Electron Chirality in Chiral Molecules (ECCM) [11].
-The value of ECCM is not large, and phenomena related to ECCM are not so many.
-Nevertheless, some phenomena are studied by our group.
+The value of ECCM is small, and phenomena related to ECCM are not extensively documented.
+Nevertheless, our research group has investigated several associated phenomena.
 One topic is the homochirality in nature [12].
-In an enantiomeric pair, the number of the left-handed electron is different, though the total electron number is the same.
-Hence, one enantiomer has a larger reaction rate of weak interaction than the other one.
-As a result, one type of an enantiomeric pair is more lost by collisions with particles with weak charge in space or on the earth [6].
-Another example is the Chirality-Induced Spin Selectivity (CISS) [13].
-CISS is the strong spin polarization when electrons pass through chiral organic molecules.
-It has not been clarified what makes large spin polarization.
-For this puzzle, the study from the viewpoint of the torque may be a key to the solution instead of the spin-orbit interaction, which is in the concept of energy.
-The torque for the electron spin is dependent on the electron chirality density, and the torques induced by the left- and right-handed electrons are opposite.
-ECCM predicts opposite electron chirality distribution in enantiomeric pair molecules.
-The zeta force may be the driving torque for the CISS.
-The correlation between the spin polarization of the CISS and the magnitude of the zeta force was confirmed for alanine and helicene molecules [14].
+In an enantiomeric pair, the number of the left-handed electron differs, while the total number of electrons is the same.
+Consequently, one enantiomer exhibits a higher reaction rate for weak interactions than its counterpart.
+This disparity leads to the preferential loss of one enantiomer through collisions with weakly charged particles in both terrestrial and extraterrestrial environments [6].
+Another noteworthy phenomenon is Chirality-Induced Spin Selectivity (CISS) [13].
+CISS is the strong spin polarization when electrons traverse chiral organic molecules such as DNA.
+The mechanisms underlying this significant spin polarization remain unclear.
+For this puzzle, examining the problem from the perspective of torque may provide critical insights, as opposed to focusing solely on spin-orbit interactions, which are typically framed in terms of energy.
+The torque experienced by electron spin is influenced by electron chirality density, with the torques generated by left- and right-handed electrons being oppositional.
+ECCM predicts an opposite distribution of electron chirality within enantiomeric pair molecules, suggesting that the zeta force may serve as the driving torque for CISS.
+The correlation between spin polarization in CISS and the magnitude of the zeta force was confirmed for alanine and helicene molecules [14].
 […]
-In the present paper, it is confirmed that the enhancement by electronic excitation also occurs for the MEC.
-As target molecules, H2Te2 and H2Se2 are chosen.
+In the present paper, it is confirmed that electronic excitation also leads to an enhancement of MEC.
+Target molecules chosen for this investigation are H2Te2 and H2Se2.
+These H2X2 type molecules are often employed to study parity-violating effects in molecules [9, 17, 20].
 In Refs.
 [7, 8], H2O2 and H2S2 are also studied for the enhancement of the PVED.
-For the MEC, the computation of molecules including only light elements with small spin-orbit interaction such as H2O2 and H2S2 requires very large basis sets and high cost [6, 11].
-Hence, H2O2 and H2S2 are omitted in the present study for the excited states.
-However, we consider that H2O2 and H2S2 will show results similar to H2Te2 and H2Se2 due to the similar patterns of the total chirality in the ground state [6, 11] and the enhancement of the PVED [7, 8].
-After the enhancement of the MEC is confirmed, it is studied whether the CBE hypothesis can also be applicable to the enhancement of the MEC.
+However, computing MEC for molecules that consist solely of light elements with small spin-orbit interactions, such as H2O2 and H2S2 necessitates large basis sets and incurs significant computational costs [6, 11].
+Specifically, in Ref.
+[6], it was demonstrated that the dyall.ae3z basis set is insufficient for accurately calculating the MEC of H2S2, while the dyall.ae4z basis set is required for reliable results.
+Moreover, for the calculation of the MEC of H2O2, even the dyall.ae4z basis set is insufficient.
+The accurate calculation of H2O2 requires additional diffuse functions [11].
+Consequently, H2O2 and H2S2 are excluded from the present study focused on excited states.
+Nonetheless, we anticipate that H2O2 and H2S2 will exhibit results analogous to those of H2Te2 and H2Se2 due to the similar patterns of the total chirality in the ground state [6, 11] and the enhancement of the PVED [7, 8].
+Following the confirmation of MEC enhancement, it is studied whether the CBE hypothesis is applicable to the enhancement of the MEC.
 This paper is organized as follows.
@@
 Here $E_{\rm PV,HOMO}$ is the contribution from one electron in the Kramers pair of the HOMO to $E_{\rm PV}$.
-(Two electrons in the same Kramers pair give the same contribution to the electron chirality.) As a more improved estimate, contributions from other orbitals are taken into account,
-$E_{\rm PV}({\rm IE})=\sum_{i}^{\rm unocc}r_{i}^{U}E_{\rm PV,\it i}-\sum_{i}^{\rm occ}r_{i}^{O}E_{\rm PV,\it i},$ (7)
-where $E_{\rm PV,\it i}$ is the contribution from one electron in the $i$-th occupied or unoccupied orbital, and $r_{i}^{O}$ ($r_{i}^{U}$) is the excitation ratio of the electron in the $i$-th occupied orbital (to the $i$-th unoccupied orbital).
-The former simple estimate is quite useful, since only Hartree-Fock (HF) or density functional theory computations are needed, while the improved estimate (IE) requires a computation beyond self-consistent field computations.
+(Two electrons in the same Kramers pair give the same contribution to the electron chirality.)
 While this CBE hypothesis is about the enhancement of the PVED, the enhancement of the MEC is considered to occur in the same cancellation-breaking mechanism.
@@
 $-\int d^{3}x\psi_{\rm HOMO}^{\dagger}(x)\gamma_{5}\psi_{\rm HOMO}(x),$ (7)
-where $\psi_{\rm HOMO}$ is the wave function of the HOMO. The improved estimate with other orbital contributions is given in the following expression,
-$\sum_{i}^{\rm unocc}r_{i}^{U}\int d^{3}x\psi_{i}^{\dagger}(x)\gamma_{5}\psi_{i}(x)-\sum_{i}^{\rm occ}r_{i}^{O}\int d^{3}x\psi_{i}^{\dagger}(x)\gamma_{5}\psi_{i}(x),$ (9)
-where $\psi_{i}$ is the wave function of one of the Kramers pair of the $i$-th orbital. In this paper, the CBE hypothesis is checked for the MEC.
+where $\psi_{\rm HOMO}$ is the wave function of the HOMO. In this paper, the CBE hypothesis is checked for the MEC.
 III Computational detail
@@
 The enhancement of the MEC as well as the PVED is confirmed.
-While the MEC in the ground state on the dihedral angle is a sigmoidal curve, in excited states a different pattern of the dependence is seen.
+While the MEC dependence on the dihedral angle in the ground state is a sigmoidal curve, in excited states a different pattern of the dependence is seen.
 All curves of the MEC have similar distributions.
 In their patterns, the MEC increases to 90 and flips the sign around $\phi=90^{\circ}$.
+The structure at $\phi=90^{\circ}$ is chiral, while the MEC is zero for the GS and some excited states.
+The reason why the PVED is zero around $\phi=90^{\circ}$ in the ground state is explained in terms of the parity violation operator $\sigma\cdot p$ and spin-orbit interaction using perturbation theory in Ref.
+[33].
+Since the MEC and PVED has the same dependence on $\phi$, the same discussion will clarify why the MEC is zero around $\phi=90^{\circ}$ in the ground state.
 Most excited states are derived as the excitation of one electron in the HOMO in the ground state.
@@
 In Fig.
-3, the estimate of the MEC in the CBE hypothesis is shown as well as the improved estimate as introduced in Sec.
-II.
+3, the estimate of the MEC in the CBE hypothesis is shown.
 The predictions by the CBE hypothesis well realize the patterns of the MEC in 1a and 1b states.
@@
 The CBE hypothesis is considered to apply the MEC.
-At most points, the simple CBE estimate does not deviate from the prediction by IE.
-This is because the electron in the HOMO is dominantly excited at most points.
-The prediction by the improved estimate seems to be worse than the simple CBE estimate around $\phi=90^{\circ}$.
-This is because there are few data for IE points.
-If there were additional data around $\phi=90^{\circ}$, the IE prediction was better than Fig.
-3.
 (a) H2Te2